# Fast Crystal Tensor Property Prediction: A General O(3)-Equivariant Framework Based on Polar Decomposition

### v2

@@ Abstract
Predicting the tensor properties of crystalline materials is a fundamental task in materials science. Unlike single-value property prediction, which is inherently invariant, tensor property prediction requires maintaining $O(3)$ group tensor equivariance. This equivariance constraint often introduces tremendous computational costs, necessitating specialized designs for effective and efficient predictions. To address this limitation, we propose a general $O(3)$-equivariant framework for fast crystal tensor property prediction, called GoeCTP. Our framework is efficient as it does not need to impose equivalence constraints onto the network architecture. Instead, GoeCTP captures the tensor equivariance with a simple external rotation and reflection (R&R) module based on polar decomposition. The crafted external R&R module can rotate and reflect the crystal into an invariant standardized crystal position in space without introducing extra computational cost. We show that GoeCTP is general as it is a plug-and-play module that can be smoothly integrated with any existing single-value property prediction framework for predicting tensor properties. Experimental results indicate that GoeCTP achieves higher prediction performance and runs 13$\times$ faster compared to existing state-of-the-art methods in elastic benchmarking datasets, underscoring its effectiveness and efficiency.
@@ 1 Introduction
The tensor properties of crystalline materials can capture intricate material responses through high-order tensors, with wide-ranging applications in fields such as physics, electronics, and engineering (Yan et al., 2024b). Compared to single-value property prediction, predicting tensor properties of crystalline materials is substantially more complex. This complexity arises from the fact that tensor properties describe how crystals respond to external physical fields, such as electric fields or mechanical stress (Nye, 1985; Resta, 1994; Yan et al., 2024b). Consequently, tensor property prediction modeling necessitates preserving consistency with the crystal’s spatial position, exhibiting “special” $O(3)$ equivariance (Yan et al., 2024b). Furthermore, high-order tensor property prediction is computationally intensive due to its high dimensionality and large data volume. Therefore, providing fast and accurate predictions of tensor properties across various materials is challenging.
Thus far, several works have been dedicated to crystal tensor property prediction. One prominent category of approaches involves ab initio physical simulation techniques, such as density functional theory (DFT) (Petousis et al., 2016). These classical simulation techniques can accurately calculate various material properties, including electronic structures, phonon spectra, and tensor properties. However, they necessitate extensive computational resources due to the complexity of handling a vast number of atoms and electrons in crystal systems (Yan et al., 2024b), hindering their applicability in practice. Alternatively, machine learning (ML) models have been proposed to facilitate the process of crystalline material property prediction. These methods typically leverage high-precision datasets deriving from ab initio simulations and utilize crystal graph construction techniques along with graph neural networks (GNNs) (Chen et al., 2019; Louis et al., 2020; Choudhary & DeCost, 2021; Xie & Grossman, 2018) or transformers (Yan et al., 2024a; Taniai et al., 2024; Yan et al., 2022; Lee et al., 2024; Wang et al., 2024a).
@@
Existing ML methods have primarily focused on predicting single-value properties of crystals. They exhibit promising prediction performance by enforcing the periodic invariance constraint into the model design. However, these methods cannot be directly applied to tensor property prediction as they overlook the inherent anisotropy of the crystal systems, failing to achieve “special” $O(3)$ equivariance (Yan et al., 2024a; Taniai et al., 2024; Yan et al., 2022; Wang et al., 2024a). Some recent works attempt to ensure equivariance through specialized designs of network architectures (Mao et al., 2024; Lou & Ganose, 2024; Heilman et al., 2024; Wen et al., 2024; Pakornchote et al., 2023; Yan et al., 2024b; Zhong et al., 2023). These methods generally employ harmonic decomposition to achieve equivariance, where the tensor space is decomposed into the direct sum of irreducible representations of the rotation group. In this regard, numerous computationally intensive operations, such as tensor products and the merging of irreducible representations, are required. These operations incur large amounts of computation overhead, particularly when processing high-order data.
To address the above challenges, we propose a novel $O(3)$-equivariant framework for fast crystal tensor property prediction dubbed GoeCTP. In particular, instead of enforcing the equivariance into the model design, we craft a simple yet effective “rotation and reflection” (R&R) module based on polar decomposition. Our designed R&R module rotates and reflects the input crystal with different positions into a standardized invariant position in space. This standardized crystal position is subsequently passed into the property prediction network to obtain an invariant tensor value. Meanwhile, the orthogonal matrix obtained from the R&R module is used to achieve equivariant tensor properties prediction. Our approach is plug-and-play as it can be readily integrated with any existing single-value property prediction network for predicting tensor properties without incurring additional computational costs. Compared to the previous state-of-the-art work Yan et al. (2024b), the GoeCTP method achieves higher quality prediction results and runs more than 13$\times$ faster in the elastic benchmarking dataset.
2 Preliminaries and problem statement
@@ 2.1 Preliminaries
The structure of crystalline materials consists of a periodic arrangement of atoms in 3D space, with a repeating unit called a unit cell. An entire crystal typically can be characterized by describing the parameters of a single unit cell, such as the types and coordinates of the atoms within it, as well as the lattice parameters (Yan et al., 2022; Wang et al., 2024b; Jiao et al., 2024). General methods for describing crystals can be divided into the Cartesian coordinate system and the fractional coordinate system, as described below.
Cartesian Coordinate System. A crystal can be mathematically represented as $\mathbf{M}=(\mathbf{A},\mathbf{X},\mathbf{L})$, where $\mathbf{A}=[\bm{a}_{1},\bm{a}_{2},\cdots,\bm{a}_{n}]^{T}\in\mathbb{R}^{n\times d_{a}}$ denotes atom features for $n$ atoms within a unit cell. Each $\bm{a}_{i}\in\mathbb{R}^{d_{a}}$ is a $d_{a}$-dimensional feature vector characterizing an individual atom. The matrix $\mathbf{X}=[\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}]^{T}\in\mathbb{R}^{n\times 3}$ contains the 3D Cartesian coordinates of $n$ atoms in the unit cell. The lattice maxtrix $\mathbf{L}=[\bm{l}_{1},\bm{l}_{2},\bm{l}_{3}]\in\mathbb{R}^{3\times 3}$ consists of the lattice vectors $\bm{l}_{1}$, $\bm{l}_{2}$, and $\bm{l}_{3}$, which form the basis of the 3D Euclidean space. A complete crystal is therefore represented as $(\hat{\mathbf{A}},\hat{\mathbf{X}})=\{(\hat{\bm{a}_{i}},\hat{\bm{x}}_{i})|\hat{\bm{x}}_{i}=\bm{x}_{i}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3},\hat{\bm{a}_{i}}=\bm{a}_{i},k_{1},k_{2},k_{3}\in\mathbb{Z},i\in\mathbb{Z},1\leq i\leq n\}$. In this representation, the integers $k_{i}$ and $l_{i}$ denote all possible atomic positions in the periodic lattice.
Fractional Coordinate System. Instead of using the standard orthogonal basis, fractional coordinate system utilizes the lattice matrix $\mathbf{L}=[\bm{l}_{1},\bm{l}_{2},\bm{l}_{3}]\in\mathbb{R}^{3\times 3}$ as the basis vectors for atomic positions. With this representation, the position of an atom is given by a fractional coordinate vector denoted as $\bm{f_{i}}=[f_{1},f_{2},f_{3}]^{T}\in[0,1)^{3}$. The corresponding Cartesian coordinate vector can then be expressed as $\bm{x}_{i}=f_{i}\bm{l}_{i}$. Therefore, for a crystal $\mathbf{M}$, it can be represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, where $\mathbf{F}=[\bm{f}_{1},\cdots,\bm{f}_{n}]^{T}\in[0,1)^{n\times 3}$ represents the fractional coordinates of all atoms in the unit cell.
Figure 1: The illustration of $O(3)$-equivariance for crystal tensor prediction. Some of the visualizations in the figure are generated using the VESTA (Momma & Izumi, 2011), Yan et al. (2024b), and VELAS (Ran et al., 2023).
@@ 2.2 Problem statement
Crystal Tensor Prediction. The crystal tensor properties prediction is a classic regression task. Its goal is to estimate the high-order tensor property denoted as $\mathcal{Y}_{label}$ from the raw crystal data represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$ while minimizing the expected error between the actual property $\mathcal{Y}_{label}$ and the estimated property value $\mathcal{Y}_{pred}$. This problem can be formally formulated as follows:
$\mathop{\min}\limits_{\theta}\sum\limits_{n=1}^{N}{||{\cal Y}_{pred}^{(n)}-{\cal Y}_{label}^{(n)}|{|^{2}}},\quad{\cal Y}_{pred}^{(n)}={f_{\theta}}(\mathbf{A},\mathbf{F},\mathbf{L}),$ (1)
@@
where $f_{\theta}(\cdot)$ represents a tensor prediction model with learnable parameters $\theta$, and the superscript $n$ denotes individual samples in the dataset. In what follows, we will omit superscript $n$ for simplicity. As did in the literature (Mao et al., 2024; Wen et al., 2024; Yan et al., 2024b), our objective is to estimate high-order tensor properties, including dielectric tensor (i.e., $\mathcal{Y}_{label}\buildrel\Delta\over{=}\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$) and elastic tensor (i.e., $\mathcal{Y}_{label}\buildrel\Delta\over{=}C\in\mathbb{R}^{3\times 3\times 3\times 3}$), respectively.
@@
$\mathbf{O(3)}$-Equivariance. The $O(3)$ group, also known as the orthogonal group in 3D, consists of all rotations and reflections in 3D space, i.e., $3\times 3$ orthogonal matrices. In predicting the crystal tensor properties, the requirements for $O(3)$ equivariance typically differ from the $O(3)$-equivariance defined in the general molecular studies (Hoogeboom et al., 2022; Song et al., 2024). Specifically, taking the dielectric tensor as an example where $\mathcal{Y}_{label}\buildrel\Delta\over{=}\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$, for a tensor prediction model $f_{\theta}(\cdot)$ in Eq. 1, if it is $O(3)$ equivariant, it must satisfy the following equality formulated as:
$f_{\theta}(\mathbf{A},\mathbf{F},\mathbf{Q}\mathbf{L})=\mathbf{Q}f_{\theta}(\mathbf{A},\mathbf{F},\mathbf{L})\mathbf{Q}^{T},$ (2)
@@
where $\mathbf{Q}\in\mathbb{R}^{n\times n}$ is an arbitrary orthogonal matrix (Yan et al., 2024b). For clarity, an illustration of $O(3)$-equivariance for crystal tensor prediction is shown in Fig. 1.
Our Objective. Our objective is to devise a new framework that can capture the $O(3)$-equivariance for accurately predicting the high-order crystal tensor properties.
@@
Figure 2: The Illustration of GoeCTP. To begin with, (1) the R&R module rotates and reflects the input crystal structure, which may have an arbitrary direction, to the standardized crystal position. Next, (2) Crystal Graph Construction module organizes the adjusted input into a crystal graph, followed by (3) the Node & Edge Feature Embedding module, which encodes the features of the crystal graph. Subsequently, (4) the Prediction module leverages these embedded features to predict the invariant tensor properties corresponding to the standardized crystal position. Finally, (5) the Reverse R&R module applies the orthogonal matrix $\mathbf{Q}$, obtained from the R&R module, to ensure the output of the equivariant tensor properties.
@@ 3 Methodology
3.1 The Rationale of Our Framework
To enforce $O(3)$ equivariance, as described in Eq. 2, existing crystal tensor prediction methods typically employ specialized network architectures based on harmonic decomposition, resulting in substantial computational overhead. For example, existing work (Yan et al., 2024b; a) gather the rotational information from neighboring nodes to the central node $i$ through the following computation formulated:

### v3

# Revisiting the Canonicalization for Fast and Accurate Crystal Tensor Property Prediction

@@ Abstract
Predicting the tensor properties of crystalline materials is a fundamental task in materials science. Unlike single-value property prediction, which is inherently invariant, tensor property prediction requires maintaining $O(3)$ group tensor equivariance. Such equivariance constraint often requires specialized architecture designs to achieve effective predictions, inevitably introducing tremendous computational costs. Canonicalization, a classical technique for geometry, has recently been explored for efficient learning with symmetry. In this work, we revisit the problem of crystal tensor property prediction through the lens of canonicalization. Specifically, we demonstrate how polar decomposition, a simple yet efficient algebraic method, can serve as a form of canonicalization and be leveraged to ensure equivariant tensor property prediction. Building upon this insight, we propose a general $O(3)$-equivariant framework for fast and accurate crystal tensor property prediction, referred to as GoeCTP. By utilizing canonicalization, GoeCTP achieves high efficiency without requiring the explicit incorporation of equivariance constraints into the network architecture. Experimental results indicate that GoeCTP achieves the best prediction accuracy and runs up to 13$\times$ faster compared to existing state-of-the-art methods in benchmarking datasets, underscoring its effectiveness and efficiency.
@@ 1 Introduction
The tensor properties of crystalline materials can capture intricate material responses through high-order tensors, with wide-ranging applications in fields such as physics, electronics, and engineering [51]. These tensor properties span various orders, such as dielectric tensor with two orders, piezoelectric tensor with three orders, and elastic tensor with four orders. Accurate prediction of these tensor properties is critical for novel materials discovery and design with targeted characteristics. Thus far, several works have been dedicated to crystal tensor property prediction. One prominent category involves ab initio physical simulation techniques, such as density functional theory (DFT) [35]. These classical simulation techniques can provide acceptable error margin for predicting various material properties. However, they necessitate extensive computational resources due to the complexity of handling large crystals with a vast number of atoms and electrons [51], hindering their applicability in practice.
Alternatively, machine learning (ML) models have been proposed to facilitate the process of crystalline material property prediction. These methods typically leverage high-precision datasets deriving from ab initio simulations and utilize crystal graph construction techniques along with graph neural networks (GNNs) [1, 26, 2, 47] or transformers [50, 42, 49, 20, 44]. Most existing methods are designed for scalar property prediction, focusing on achieving $SO(3)$ invariance of the crystal structures. However, these scalar-property methods could not be adapted to predict tensor properties of crystalline materials, which exhibits significantly higher complexity. This complexity arises from the fact that tensor properties describe how crystals respond to external physical fields, such as electric fields or mechanical stress [33, 38, 51]; their modeling necessitates preserving consistency with the crystal’s spatial orientation, exhibiting a unique tensor $O(3)$ equivariance [51].
@@
Therefore, a few recent studies attempt to ensure equivariance through specialized designs of network architectures [28, 25, 11, 46, 34, 51, 53]. These methods generally employ harmonic decomposition to achieve equivariance, where the tensor space is decomposed into the direct sum of irreducible representations of the rotation group. In this regard, numerous computationally intensive operations, such as tensor products and the merging of irreducible representations, are required. These operations incur large amounts of computation overhead, particularly when processing high-order data. Therefore, providing fast and accurate predictions of tensor properties across various materials is challenging.
One approach that has show promising results in efficient learning with symmetry across various fields, including point clouds [22], n-body simulation [17], and antibodies generation [29], is the use of “canonicalization". Specifically, canonicalization maps a geometric data to an invariant representation [27, 6], referred to as the canonical form, and subsequently reconstructs an equivariant output from the canonical form without imposing any architectural constraints on the network. However, to date, it has not been explored for crystal tensor property prediction.
In this work, we revisit the task of crystal tensor property prediction through the lens of canonicalization and introduce a novel canonicalization strategy tailored for this particular setting. Specifically, rather than embedding equivariance directly into model architecture, we propose a simple yet effective canonicalization module, termed R&R. In particular, our canonicalization is instantiated based on polar decomposition, a continuous mapping technique [6] that can provide enhanced robustness. During prediction, the R&R module applies rotations and reflections to transform the input crystal structure into its canonical form. The canonical form is then fed into a property prediction network to obtain the corresponding canonical tensor representation. Simultaneously, the orthogonal matrix derived from the R&R module is utilized to recover the equivariant output via the tensor transformation rule, enabling equivariant tensor property prediction without incurring additional computational overhead. We conducted experiments on dielectric, piezoelectric, and elastic tensor datasets, respectively, showcasing that the proposed method can achieve $O(3)$-equivariant tensor prediction while maintaining high efficiency. Compared to the previous state-of-the-art work Yan et al. [51], the GoeCTP method achieves higher prediction accuracy and runs by up to 13$\times$ faster in benchmarking datasets.
2 Background
@@ 2.1 Preliminaries
The structure of crystalline materials consists of a periodic arrangement of atoms in 3D space, with a repeating unit called a unit cell. An entire crystal typically can be characterized by describing the parameters of a single unit cell, such as the types and coordinates of the atoms within it, as well as the lattice parameters [49, 45, 16]. General methods for describing crystals can be divided into the Cartesian coordinate system and the fractional coordinate system, as described below.
Cartesian Coordinate System. A crystal can be mathematically represented as $\mathbf{M}=(\mathbf{A},\mathbf{X},\mathbf{L})$, where $\mathbf{A}=[\bm{a}_{1},\bm{a}_{2},\cdots,\bm{a}_{n}]^{\top}\in\mathbb{R}^{n\times d_{a}}$ denotes atom features for $n$ atoms within a unit cell. Each $\bm{a}_{i}\in\mathbb{R}^{d_{a}}$ is a $d_{a}$-dimensional feature vector characterizing an individual atom. The matrix $\mathbf{X}=[\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}]^{\top}\in\mathbb{R}^{n\times 3}$ contains the 3D Cartesian coordinates of $n$ atoms in the unit cell. The lattice maxtrix $\mathbf{L}=[\bm{l}_{1},\bm{l}_{2},\bm{l}_{3}]\in\mathbb{R}^{3\times 3}$ consists of the lattice vectors $\bm{l}_{1}$, $\bm{l}_{2}$, and $\bm{l}_{3}$, which form the basis of the 3D Euclidean space. A complete crystal is therefore represented as $(\hat{\mathbf{A}},\hat{\mathbf{X}})=\{(\hat{\bm{a}_{i}},\hat{\bm{x}}_{i})|\hat{\bm{x}}_{i}=\bm{x}_{i}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3},\hat{\bm{a}_{i}}=\bm{a}_{i},k_{1},k_{2},k_{3}\in\mathbb{Z},i\in\mathbb{Z},1\leq i\leq n\}$. In this representation, the integers $k_{i}$ and $l_{i}$ denote all possible atomic positions in the periodic lattice.
Fractional Coordinate System. Instead of using the standard orthogonal basis, fractional coordinate system utilizes the lattice matrix $\mathbf{L}=[\bm{l}_{1},\bm{l}_{2},\bm{l}_{3}]\in\mathbb{R}^{3\times 3}$ as the basis vectors for atomic positions. With this representation, the position of an atom is given by a fractional coordinate vector denoted as $\bm{f_{i}}=[f_{1},f_{2},f_{3}]^{\top}\in[0,1)^{3}$. The corresponding Cartesian coordinate vector can then be expressed as $\bm{x}_{i}=f_{i}\bm{l}_{i}$. Therefore, for a crystal $\mathbf{M}$, it can be represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, where $\mathbf{F}=[\bm{f}_{1},\cdots,\bm{f}_{n}]^{\top}\in[0,1)^{n\times 3}$ represents the fractional coordinates of all atoms in the unit cell.
Figure 1: The illustration of $O(3)$-equivariance for crystal tensor prediction. Some of the visualizations in the figure are generated using the VESTA [30], Yan et al. [51], and VELAS [37].
@@ 2.2 Problem statement
Crystal Tensor Prediction. The crystal tensor properties prediction is a classic regression task. Its goal is to estimate the high-order tensor property denoted as $\mathcal{Y}$ from the raw crystal data represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$. Given that the crystal data $\mathbf{M}$ resides within the input space $\mathcal{V}$, and the tensor property $\mathcal{Y}$ belongs to the separate output space $\mathcal{W}$, the objective of crystal tensor prediction is to find a function ${f_{\theta}}:\mathcal{V}\to\mathcal{W}$ that accurately maps input crystal data to the desired tensor property. This is achieved by minimizing the discrepancy between the true property $\mathcal{Y}$ and the predicted property value $\mathcal{Y}_{pred}$. Therefore, the crystal tensor property prediction task can be mathematically formulated as the following optimization problem:
$\mathop{\min}\limits_{\theta}\sum\limits_{n=1}^{N}{||{\cal Y}_{pred}^{(n)}-{\cal Y}^{(n)}|{|^{2}}},\quad{\cal Y}_{pred}={f_{\theta}}(\mathbf{A},\mathbf{F},\mathbf{L}),$ (1)
@@
where $f_{\theta}(\cdot)$ represents a tensor prediction model with learnable parameters $\theta$, and the superscript $n$ denotes individual samples in the dataset. In what follows, we will omit superscript $n$ for simplicity. As did in the literature [28, 46, 51], our objective is to estimate high-order tensor properties, including 2-order dielectric tensor (i.e., $\mathcal{Y}\buildrel\Delta\over{=}\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$), 3-order piezoelectric tensor (i.e., $\mathcal{Y}\buildrel\Delta\over{=}\mathbf{e}\in\mathbb{R}^{3\times 3\times 3}$), and 4-order elastic tensor (i.e., $\mathcal{Y}\buildrel\Delta\over{=}C\in\mathbb{R}^{3\times 3\times 3\times 3}$), respectively.
$\mathbf{O(3)}$ group. The $O(3)$ group is an orthogonal group, consisting of rotations and reflections. The elements $g\in O(3)$ act on vectors or tensors in $\mathcal{V}$, $\mathcal{W}$ through their respective group representations, $\rho_{\mathcal{V}}:O(3)\to GL(\mathcal{V})$ and $\rho_{\mathcal{W}}:O(3)\to GL(\mathcal{W})$, where $GL(\mathcal{V})$ and $GL(\mathcal{W})$ are the space of invertible linear maps $\mathcal{V}\to\mathcal{V}$ and $\mathcal{W}\to\mathcal{W}$, respectively. Specifically, the action of an element $g\in O(3)$ on crystal data $\mathbf{M}$ is defined as: $g\cdot\mathbf{M}=\rho_{\mathcal{V}}(g)\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{Q}\mathbf{L})$. For dielectric tensor property $\mathcal{Y}$, the action of $g$ is given by: $g\cdot\mathcal{Y}=\rho_{\mathcal{W}}(g)\mathcal{Y}=\mathbf{Q}\mathcal{Y}\mathbf{Q}^{\top}$, where $\mathbf{Q}$ is a $3\times 3$ orthogonal matrix (for additional details regarding the transformation of tensor properties, please refer to Appendix A.8).
@@
$\mathbf{O(3)}$-Equivariance for Crystal Tensor Prediction. In the crystal tensor properties prediction task, due to the difference between $\rho_{\mathcal{V}}(g)$ and $\rho_{\mathcal{W}}(g)$, the requirements for $O(3)$ equivariance typically differ from the $O(3)$-equivariance defined in the general molecular studies [13, 48, 52, 40]. Specifically, taking the dielectric tensor as an example where $\mathcal{Y}_{label}\buildrel\Delta\over{=}\bm{\varepsilon}\in\mathbb{R}^{3\times 3}$, for a tensor prediction model $f_{\theta}(\cdot)$ in Eq. 1, if it is $O(3)$ equivariant, it must satisfy the following equality formulated as:
$f_{\theta}(\mathbf{A},\mathbf{F},\mathbf{Q}\mathbf{L})=\mathbf{Q}f_{\theta}(\mathbf{A},\mathbf{F},\mathbf{L})\mathbf{Q}^{\top},$ (2)
@@
where $\mathbf{Q}\in\mathbb{R}^{n\times n}$ is an arbitrary orthogonal matrix [51]. For clarity, an illustration of $O(3)$-equivariance for crystal tensor prediction is shown in Fig. 1 (for more equivariance details, see Appendix A.8).
@@
Figure 2: The Illustration of GoeCTP. To begin with, (1) the R&R module rotates and reflects the input crystal structure, which may have an arbitrary orientation, to the canonical form of this crystal. Next, (2) Crystal Graph Construction module organizes the adjusted input into a crystal graph, followed by (3) the Node & Edge Feature Embedding module, which encodes the features of the crystal graph. Subsequently, (4) the Prediction module leverages these embedded features to predict the canonical form of tensor properties corresponding to the canonical form of input crystal. Finally, (5) the Reverse R&R module applies the orthogonal matrix $\mathbf{Q}$, obtained from the R&R module, to ensure the equivariance of the output tensor properties.
@@ 3 Methodology
In this section, we will first provide the motivation driving our new framework. We then reformulate the $O(3)$ equivariant tensor prediction task from the perspective of canonicalization. Building upon this foundational perspective, we demonstrate how polar decomposition can be employed as a canonicalization strategy and effectively applied to solve tensor prediction tasks with both effectiveness and efficiency.
3.1 The Motivation of Our Framework
To enforce $O(3)$ equivariance, as described in Eq. 2, existing crystal tensor prediction methods typically employ specialized network architectures based on harmonic decomposition, resulting in substantial computational overhead. For example, existing work gather the rotational information from the neighboring nodes to the central node $i$ through the following computation formulated [51]: